write a general formula to describe the variation. y varies directly with x; y=5 when x=30
Answer by Alan3354(69443)

write a general formula to describe the variation. y varies directly with x; y=5 when x=30
-------------
y = kx
---
5 = k*30
k = 1/6
--> y = x/6